OpenStudy (anonymous):

The probability that Juan hits the target is 0.7. If Juan got three tries to hit the target, what is the probability that he will hit the target only in the first two tries? Find probability of hitting each time is same.

OpenStudy (cwrw238):

these probabilities (0.7) are independent so for a series of attempts they are multiplued P(hitting first and missing the third) = 0.7 * 0.7 * 0.3 note probability of missing = 1 - 0.7 = 0.3

OpenStudy (anonymous):

oh, thanks :')

OpenStudy (cwrw238):

yw